Trying to prove this is an identity.
cos^3 J sin^2 J = cos^3 J - cos^5 J

Answer by Aswathy(23)

LHS=cos^3J sin^2J
=cos^3J (1-cos^2J) (by using the identity sin^2 J+cos^2 J=1)
=cos^3 J-cos^5 J=RHS